OpenStudy (rational):

f(3x) = x + f(3x-3) f(3) =1 Find f(999)

OpenStudy (domebotnos):

f is 999??

OpenStudy (domebotnos):

oh wouldnt f be 1??

OpenStudy (rational):

f is a function here..

OpenStudy (domebotnos):

\[f(3x)=x =+ (3x-3)\]

Parth (parthkohli):

\[f(k) - f(k - 3) = k/3\]I'm sensing some telescoping here.

OpenStudy (rational):

I think that looks like a 3 level recurrence relation \[a_n -a_{n-3} = n/3,~~a_3 = 1\] impossible to have an unique solution without knowing \(a_1\) and \(a_2\)

Parth (parthkohli):

Or, rather,\[f(k+3) - f(k) = k/3 + 1\]So\[f(6) = 5\]\[f(9) = 8\]\[f(12) = 13\]FIBONACCI!

Parth (parthkohli):

\[f(15) = 6 + 13 = 19\]nvm

OpenStudy (anonymous):

is it 333*167 -1 ?

Parth (parthkohli):

So hold on, it's possible to do this by telescoping, yes?

OpenStudy (anonymous):

For \(x=1\) we get another initial point, \[f(3)=1+f(0)~~\implies~~f(0)=0\] How many more are needed to solve the recurrence @rational?

OpenStudy (rational):

that looks very huge but it seems you both are on right tack..

Parth (parthkohli):

\[f(3x + 3) - f(3x) = x+1\]\[f(6) - f(3) = 2\]\[f(9) - f(6) = 3\]\[f(12) - f(9) = 4\]yup

OpenStudy (rational):

hey @divu.mkr sry your answer is correct!

Parth (parthkohli):

\[\vdots \\ f(999) - f(996) = 334\]

OpenStudy (anonymous):

f(3x) = x+ f(3(x-1)) = x+ x-1 f(3(x-2)) = x+x-1 +x-2 .... +2 + f(3) = x+x-1 + x-2 .... +2 +1 x(x+1)/2

Parth (parthkohli):

Add all of them and\[f(3) + f(999) = \frac{333\cdot 334}{2}\]\[f(999) = 333\cdot 167 - 1\]

Parth (parthkohli):

Hold on. It should be -2.

Parth (parthkohli):

Because 1 is not included in the sum. It starts from 2.

OpenStudy (anonymous):

isnt it f(999) - f(3) ?

Parth (parthkohli):

Oh, that's right.

Parth (parthkohli):

Which means that there's no -1 in there?

OpenStudy (rational):

Yes both solutions are correct! Looks the sequence formed by picking third terms gives sum of first "x/3" natural numbers

OpenStudy (anonymous):

yep @ParthKohli

Parth (parthkohli):

@rational which is why there can't be any -1, ahhhhh.

OpenStudy (rational):

f(3x) = x(x+1)/2 looks good to me plugin x=333 to get f(3*333) = 55611 gotcha!

Parth (parthkohli):

That was simple. I wonder why we couldn't notice that in the recurrence relation itself. >_<

OpenStudy (kainui):

I found an alternate way to the recurrence relation that's kinda cute. Let's make a new function g(x)=3x and then we can make another new function h(x)=f(g(x)). Then we can transform our condition f(3)=1 into f(g(1))=h(1)=1 and the thing to find f(g(333))=h(333). Our equation then becomes \[\Large h(x)=x+h(x-1)\] substituting in x+1 we get \[\Large h(x+1)=x+1+h(x)\] and the rest is just the same technique as you would guess it to be.

OpenStudy (rational):

Thats same as divu.mkr's solution except it looks more neat :)